If Leah is 6 years older than her sister, Sue, and John is 5 years older than Leah, and the total of their ages is 41. Then how old is Sue?
Answer by ankor@dixie-net.com(12691)

Write an equation from each statement:
:
If Leah is 6 years older than her sister, Sue.
L = S + 6
:
John is 5 years older than Leah,
J = L + 5
Substitute (S+6) for L and you have:
J = (S+6) + 5
J = S + 11
:
the total of their ages is 41.
L + S + J = 41
:
Then how old is Sue?
Substitute (S+5) for L and (S+11) for J in the above equation:
(S+6) + S + (S+11) = 41
3S + 17 = 41
3S = 41 - 17
3S = 24
S = 24/3
S = 8 yrs is Sue's age
:
:
Check:
L = 8 + 6 = 14
J = 8 + 11 = 19
:
L + S + J = 41
14 + 8 + 19 = 41
:
Did this make sense to you? Any questions?